# Can Contrastive Learning Improve Class-Imbalanced Diffusion Model?

### v1

@@ 3.2 Positive Pairs and MSE loss
Conditinal-unconditinal Alignment Facilitates Knowledge Sharing for Class-imbalanced GAN We take inspiration from UTLO (utlo, ), which addresses long-tailed generation with GANs, and observes that the similarity between head class and tail class images increases at lower resolutions. To share knowledge between the head class and tail class, UTLO (utlo, ) proposes unconditional GAN objectives for lower-resolution images and conditional GAN objectives for subsequent higher-resolution images. More specifically, the lower-resolution image/feature becomes less sensitive to class labels, and the second part of the generator generates the fine details of images. Utilizing unconditional generation for lower resolution is effective in increasing the diversity and quality of tail class images. In a different setting of training GAN with limited data, Transitional-GAN (transitioncgan, ) leverages unconditional generation in a two-staged training scheme, as unconditional generation can give a better FID than conditional generation when data is limited.
@@
Conditional-unconditinal Alignment for Class-imbalanced Diffusion Model We aim to adapt the conditional-unconditional alignment approach used in GANs (utlo, ; transitioncgan, ) to diffusion models. Our key insight is that the diffusion model denoises images recursively, and the similarity between head class images and tail class images is higher during the initial timesteps of the denoising process. Our approach is different from (utlo, ), which proposes unconditional generation for lower-resolution images. We visualize the reverse processing of unconditional generation and conditional generation with different class labels of the original DDPM, starting from the same Gaussian noise, as shown in Fig. 3 with more examples in appendix A.3.4. Images from unconditional generation and conditional generation share similar low-frequency components, e.g., the coarse shape of synthesized objects and background region, especially at the initial time steps si2024freeu .
Figure 3: The leftmost column shows synthetic images with different classes but with the same initial noise or random seed. We visualize the low-frequency component and the high-frequency component from the reverse processing. It shows that low-frequency components are similar during initial time steps for different classes with the same initial noise si2024freeu . More examples are in appendix A.3.4.
@@
Motivated by this observation, we propose to match unconditional generation and conditional generation at the initial time steps, to enable knowledge sharing between tail classes and head classes with abundant data. Another motivation is that class conditions are not necessary for all timesteps. In fact, using adaptive guidance weight with a small weight in the initial steps can yield a better FID compared to using constant guidance weight cfgweightscheduler .
@@
From Conditional-unconditinal Distributin Maching to MSE Loss We penalize KL divergence between the conditional distribution $p_{\theta}(x_{t-1}|x_{t}^{i},c^{i})$ and unconditional distribution $p_{\theta}(x_{t-1}|x_{t}^{i})$:
$\mathcal{L}^{i,t}_{mseclr}=D_{KL}[p_{\theta}(x_{t-1}|x_{t}^{i},c^{i})||p_{\theta}(x_{t-1}|x_{t}^{i})].$ (2)
@@
$\mathcal{L}^{i,t}_{mseclr}=\mathbb{E}[\frac{1}{2\sigma_{t}^{2}}||\mu_{\theta}(x_{t}^{i},t,c^{i})-\mu_{\theta}(x_{t}^{i},t)||^{2}]+C,$ (3)
@@
Then, $\mathcal{L}_{mseclr}$ simplifies to:
$\mathcal{L}^{i,t}_{mseclr}\propto||\epsilon_{\theta}(x_{t}^{i},t,c^{i})-\epsilon_{\theta}(x_{t}^{i},t)||^{2}.$ (5)
As shown in Fig. 2, the anchor vector $h_{anc}^{i}$ is fed into decoder $d(\cdot)$ to get unconditional noise estimation $\epsilon_{unc}^{i}:=\epsilon_{\theta}(x_{t}^{i};t,\emptyset)$. Meanwhile, the anchor image $x_{t}^{i}$ is also incorporated with the class label condition $\mathbf{c^{i}}$ fed into the network to get the conditional noise estimation $\epsilon_{cond}^{i}:=\epsilon_{\theta}(x_{t}^{i};t,\mathbf{c^{i}})$. The MSE contrastive loss is defined between the unconditional and conditional estimation:  [^1] Our MSE loss is similar to CBDM loss cbdm . We minimize the distance between conditional generation and unconditional generation in image space, while CBDM minimizes distances between conditional generations for two different conditions. Our motivation is different from CBDM, but we end up with a similar but different loss. Our work offers different interpretations of previous work and makes insightful connections.
$\mathcal{L}_{mseclr}=\frac{1}{|B|}\sum_{i\in B}{\mathbb{E}_{t,x_{0}}[\frac{t}{T}\|\epsilon_{\theta}(x_{t}^{i};t,\mathbf{c^{i}})-\epsilon_{\theta}(x_{t}^{i};t,\emptyset)\|^{2}}].$ (6)
@@
3.3 Overall Framework
Algorithm 1 Training algorithm of CLDM.
Set $\mathcal{L}_{ddpm},\mathcal{L}_{nceclr},\mathcal{L}_{mseclr}=0$
@@ 3.3 Overall Framework
     $\mathcal{L}_{nceclr}=\mathcal{L}_{nceclr}+(-\log\frac{\pi_{anc}}{\pi_{anc}+\pi_{neg}})$
@@
         $\mathcal{L}_{mseclr}=\mathcal{L}_{mseclr}+\frac{t}{T}\|\epsilon_{unc}^{i}-\epsilon_{cond}^{i}\|^{2}$
@@
$\mathcal{L}=\frac{1}{|B|}(\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}+\gamma\cdot\mathcal{L}_{mseclr})$
Our final loss function is the sum of the original DDPM loss and our contrastive losses:
$\mathcal{L}=\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{nceclr}+\gamma\cdot\mathcal{L}_{mseclr},$ (7)
where $\mathcal{L}_{ddpm}$ is a standard mse loss for matching the estimated noise item with the ground-truth noise item, same as in DDPM ddpm $\alpha$ and $\gamma$ are the hyper-parameters for our InfoNCE loss and MSE loss.
The overall algorithm framework is shown as Algorithm 1. In classifier-free guidance (CFG), the class labels $\mathbf{c}$ are randomly dropped by a specific probability $p_{uncond}$. In our setting, we keep the $p_{uncond}=10\%$, same as (cbdm, ). The parameters of the noise estimation network $\epsilon_{\theta}$ are updated by the gradient of the final loss $\nabla_{\theta}\mathcal{L}$ as in Eq. 7.
@@ 4.1 Experimental Setup
DatasetsWe report results on multiple datasets, including the long-tailed versions of CIFAR10/CIFAR100 (cifar, ), TinyImageNet (tinyimagenet, ), ImageNet (imagenet, ), and Places365 (places, ), namely CIFAR10LT/CIFAR100LT (cifarlt, ), TinyImageNetLT, ImageNetLT, and PlacesLT. These are commonly used datasets in long-tailed image generation (cbdm, ; diffrop, ; oclt, ) and recognition (cifarlt, ). Same as (cbdm, ; diffrop, ; oclt, ), we adopt the same methodology presented in (cifarlt, ) to construct long-tailed version datasets with an imbalance factor of $0.01$. We keep the original image size of $32\times 32$ for CIFAR10LT/CIFAR100LT, while resizing images to $64\times 64$ for TinyImageNetLT, ImageNetLT and PlacesLT.
Baselines We compare our proposed methods with multiple baselines, including: 1) original DDPM (ddpm, ), 2) CBDM (cbdm, ), which leverages imbalanced ground-truth label and balanced pseudo-labels for co-training, and 3) OCLT (oclt, ), which facilitates feature transfer from head classes to tail classes. We also compare to a widely used data augmentation technique, Adaptive Discriminator Adaption (ADA) (karras2020training, ) for generative models on the full DDPM 1000 sampling steps.
Evaluation MetricsWe use standard metrics to evaluate both the fidelity and diversity of generated samples, including Fréchet Inception Distance (FID) (fid, ), Inception Score (IS) (is, ), FID of tail classes (FIDtail), and Kernel Inception Distance (KID) (kid, ). FID, IS, and KID are regarded as overall quality metrics, while FIDtail are more inclined to measure quality for tail classes, which are widely used in evaluating generative models on long-tailed scenes (cbdm, ; oclt, ; utlo, ; xu2024rethinking, ). More implementation details are provided in the appendix A.1.
@@ 4.2 Quantitative Results
[…]
Table 1: Our method outperforms other baselines on all datasets. We also provide the results of DDPM trained on balanced datasets, which show the upper bound of performance. All models are measured with DDIM (ddim, ) 100 steps for conditional generation with CFG. red ‘()’ shows improvement of our method over DDPM baseline (ddpm, ). blue ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.
@@
DDPM (ddpm, ) 5.93 12.25 1.64
CBDM (cbdm, ) 5.81 10.01 1.58
OCLT (oclt, ) 6.10 11.13 1.58
CLDM (Ours) 5.56 (-0.37) 10.03 (-2.22) 1.27 (-0.37)
@@
DDPM (ddpm, ) 6.95 17.45 1.71
CBDM (cbdm, ) 6.50 17.36 1.41
OCLT (oclt, ) 6.45 17.22 1.42
CLDM (Ours) 6.24 (-0.71) 16.35 (-1.10) 1.36 (-0.35)
@@
DDPM (ddpm, ) 18.67 40.12 6.31
CBDM (cbdm, ) 20.90 48.07 6.55
OCLT (oclt, ) 17.72 39.67 5.61
CLDM (Ours) 17.16 (-1.51) 37.88 (-2.24) 4.89 (-1.42)
ImageNetLT DDPM (ddpm, ) 16.89 35.24 7.87
CBDM (cbdm, ) 17.38 37.22 8.04
OCLT (oclt, ) 16.45 33.82 7.04
CLDM (Ours) 16.16 (-0.73) 33.93 (-1.31) 6.60 (-1.27)
PlacesLT DDPM (ddpm, ) 13.89 23.74 5.26
CBDM (cbdm, ) 15.15 26.07 5.64
OCLT (oclt, ) 13.04 22.84 4.17
CLDM (Ours) 11.99 (-1.90) 20.84 (-2.90) 3.63 (-1.63)
Table 2: We compare our method with other baselines on CIFAR10LT/CIFAR100LT datasets, with DDPM 1000 steps for conditional generation with CFG. We also report the data augmentation method ADA karras2020training , which is orthogonal to our method.
      Method       CIFAR10LT       CIFAR100LT

### v2

@@ 3.2 Conditional-unconditional Alignment Loss
Conditinal-unconditinal Alignment Facilitates Knowledge Sharing for Class-imbalanced GAN We take inspiration from UTLO (Khorram et al., 2024), which addresses long-tailed generation with GANs, and observes that the similarity between head class and tail class images increases at lower resolution representations. To share knowledge between the head class and tail class, UTLO (Khorram et al., 2024) proposes unconditional GAN objectives for lower-resolution representations and conditional GAN objectives for subsequent higher-resolution images. More specifically, the lower-resolution representations becomes less sensitive to class labels, and the second part of the generator generates the fine details of images. Utilizing unconditional generation for lower resolution is effective in increasing the diversity and quality of tail class images. In a different setting of training GAN with limited data, Transitional-GAN (Shahbazi et al., 2022) leverages unconditional generation in a two-staged training scheme, as unconditional generation can give a better FID than conditional generation when data is limited.
@@
Conditional-unconditinal Alignment for Class-imbalanced Diffusion Model We aim to adapt the conditional-unconditional alignment approach used in GANs (Khorram et al., 2024; Shahbazi et al., 2022) to diffusion models. Our key insight is that the diffusion model denoises images recursively, and the similarity between head class images and tail class images is higher during the initial timesteps of the denoising process. This is similar but different from UTLO (Khorram et al., 2024), which proposes unconditional generation for lower-resolution representations. We visualize the reverse processing of unconditional generation and conditional generation with different class labels of the original DDPM, starting from the same Gaussian noise, as shown in Fig. 3 and more examples in appendix A.3.4. Images from unconditional generation and conditional generation share similar low-frequency components, e.g., the coarse shape of synthesized objects and background region, especially at the initial time steps (Si et al., 2024).
@@
Motivated by this observation, we propose to match unconditional generation and conditional generation at the initial time steps, to enable knowledge sharing between tail classes and head classes with abundant data. Another motivation is that class conditions are not necessary for all timesteps. In fact, using adaptive guidance weight with a small weight in the initial steps can yield a better FID compared to using constant guidance weight (Wang et al., 2024).
@@
From Conditional-unconditional Distribution Matching to Alignment Loss We penalize KL divergence between the conditional distribution $p_{\theta}(x_{t-1}|x_{t}^{i},c^{i})$ and unconditional distribution $p_{\theta}(x_{t-1}|x_{t}^{i})$:
$\mathcal{L}^{i,t}_{al}=D_{KL}[p_{\theta}(x_{t-1}|x_{t}^{i},c^{i})||p_{\theta}(x_{t-1}|x_{t}^{i})].$ (2)
@@
[…]
@@
Then, $\mathcal{L}_{al}$ simplifies to:
[…]
As shown in Fig. 2, the anchor vector $h_{anc}^{i}$ is fed into decode network d()d(*) to get unconditional noise estimation $\epsilon_{unc}^{i}:=\epsilon_{\theta}(x_{t}^{i};t,\emptyset)$. Meanwhile, the anchor image $x_{t}^{i}$ is also incorporated with the class label condition $\mathbf{c^{i}}$ fed into the network to get the conditional noise estimation $\epsilon_{cond}^{i}:=\epsilon_{\theta}(x_{t}^{i};t,\mathbf{c^{i}})$. The alignment loss is defined between the unconditional and conditional noise estimation:
$\mathcal{L}_{al}=\frac{1}{|B|}\sum_{i\in B}{\mathbb{E}_{t,x_{0}}[\frac{t}{T}\|\epsilon_{\theta}(x_{t}^{i};t,\mathbf{c^{i}})-\epsilon_{\theta}(x_{t}^{i};t,\emptyset)\|^{2}}].$ (6)
@@
Figure 3: The leftmost column shows synthetic images with different classes but with the same initial noise or random seed. We visualize the low-frequency component and the high-frequency component from the reverse processing. It shows that low-frequency components are similar during initial time steps for different classes with the same initial noise (Si et al., 2024). More examples are in appendix A.3.4.
Algorithm 1 Training algorithm of CCUA.
Set $\mathcal{L}_{ddpm},\mathcal{L}_{ucl},\mathcal{L}_{al}=0$
@@
  $\mathcal{L}_{ucl}=\mathcal{L}_{ucl}+(-\log\frac{\pi_{anc}}{\pi_{anc}+\pi_{neg}})$
@@
   $\mathcal{L}_{al}=\mathcal{L}_{al}+\frac{t}{T}\|\epsilon_{unc}^{i}-\epsilon_{cond}^{i}\|^{2}$
@@
$\mathcal{L}=\frac{1}{|B|}(\mathcal{L}_{ddpm}+\alpha\cdot\mathcal{L}_{ucl}+\gamma\cdot\mathcal{L}_{al})$
3.3 Overall Framework
Our final loss function $\mathcal{L}$ is the sum of the standard DDPM (Ho et al., 2020) loss $\mathcal{L}_{ddpm}$ and our contrastive conditional-unconditional alignment loss $\mathcal{L}_{ccua}$:
$\mathcal{L}_{ccua}=\alpha\cdot\mathcal{L}_{ucl}+\gamma\cdot\mathcal{L}_{al},$ (7)
where  $\alpha$ and $\gamma$ are the hyper-parameters for our unsupervised contrastive loss and alignment loss.
The overall algorithm framework is shown as Algorithm 1. In classifier-free guidance (CFG), the class labels $\mathbf{c}$ are randomly dropped by a specific probability $p_{uncond}$. In our setting, we keep the $p_{uncond}=10\%$, same as (Qin et al., 2023). The parameters of the noise estimation network $\epsilon_{\theta}$ are updated by the gradient of the final loss $\nabla_{\theta}\mathcal{L}$ as in Eq. 7.
@@ 4.1 Experimental Setup
We report results on multiple datasets, including the long-tailed versions of ImageNet, TinyImageNet, and Places datasets, namely ImageNet-LT, TinyImageNet-LT, and Places-LT, which are commonly used datasets in long-tailed image generation and recognition. We also report results and conduct more analysis on CIFAR10-LT/CIFAR100-LT datasets, which are shown in the appendix A.2. We measure IS Score, FID, spatial FID (Dhariwal & Nichol, 2021), Kernel Inception Distance (KID) (Bińkowski et al., 2018) as the overall quality metrics. We also report FID of tail classes (FIDtail), inclined to measure quality for tail classes. Precision and Recall are reported to respectively measure fidelity and diversity, which are widely used in evaluating generative models on long-tailed scenes (Qin et al., 2023; Zhang et al., 2024; Khorram et al., 2024). More implementation details are provided in the appendix A.1. We compare our proposed methods with multiple baselines, covering diffusion transformer and unet-based architecture. For diffusion transformer, we compare CCUA to SiT (Ma et al., 2024) and Dispersive Loss (Wang & He, 2025). For unet-based architecture, we compare to the original DDPM (Ho et al., 2020), CBDM (Qin et al., 2023), and OCLT (Zhang et al., 2024).
@@ 4.2 Quantitative Results
Table 1: Comparison on ImageNet-LT $256\times 256$ with SiT pipeline. We use Blue ‘()’ to highlight the improvement of our method over SiT baseline on FID and Recall, which denotes the overall quality and diversity, respectively.
Epochs Method IS $\uparrow$ FID $\downarrow$ sFID $\downarrow$ Prec. (%) $\uparrow$ Recall (%) $\uparrow$
40 SiT 53.93 33.87 22.66 54.56 19.17
+Dispersive Loss 53.87 34.00 22.68 54.87 19.84
+CCUA (ours) 73.60 25.89 (-7.98) 16.54 58.62 23.05 (+4.08)
80 SiT 78.88 25.72 21.95 64.35 18.58
+Dispersive Loss 83.80 25.22 21.76 65.50 17.30
+CCUA (ours) 103.66 19.99 (-5.73) 14.88 65.80 21.31 (+2.73)
120 SiT 103.11 21.20 20.17 69.34 18.25
+Dispersive Loss 104.02 21.32 20.46 68.07 18.56
+CCUA (ours) 119.11 17.55 (-3.65) 13.89 68.38 21.20 (+2.95)
160 SiT 111.70 19.95 20.12 70.34 18.61
+Dispersive Loss 115.64 19.76 20.18 69.98 18.96
+CCUA (ours) 124.87 16.41 (-3.54) 13.17 69.84 21.34 (+2.73)
Class-imbalanced Generation for Diffusion Transformer We apply the proposed CCUA loss into the standard diffusion transformer training pipeline, SiT (Ma et al., 2024), on ImageNet-LT dataset, and compare it to Dispersive Loss (Wang & He, 2025), a concurrent work of applying contrastive learning for improving SiT on balanced dataset. All models are sampled with ODE flow 50 steps for conditional generation with CFG strength 7.5. As shown in Table 1, our method achieves remarkable improvement compared to SiT and Dispersive Loss on various training epochs. Our method achieves above 10% improvement on Recall, while keep the similar Precision as SiT, illustrating the higher diversity of the proposed CCUA loss without trading off fidelity.
Table 2: Comparison on TinyImageNet-LT $64\times 64$ and Places-LT $64\times 64$ with DDPM pipeline. Blue ‘()’ shows improvement of our method over DDPM baseline. Green ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.
@@
DDPM (Ho et al., 2020) 18.67 40.12 6.31
CBDM (Qin et al., 2023) 20.90 48.07 6.55
OCLT (Zhang et al., 2024) 17.72 39.67 5.61
CCUA (ours) 15.24 (-3.43) 30.39 (-9.73) 3.83 (-2.48)
Places-LT DDPM (Ho et al., 2020) 13.89 23.74 5.26
CBDM (Qin et al., 2023) 15.15 26.07 5.64
OCLT (Zhang et al., 2024) 13.04 22.84 4.17
CCUA (ours) 11.99 (-1.90) 20.84 (-2.90) 3.63 (-1.63)
@@
Table 3: FID score for three ‘super-categories’: ‘Head’, ‘Body’, and ‘Tail’. All models are measured with DDIM (Song et al., 2020) 100 sampling steps for conditional generation with CFG.
@@
Head
$\sim 80\%$
Body
$\sim 17\%$
Tail
$\sim 3\%$
All
TinyImageNet-LT DDPM (Ho et al., 2020) 21.27 33.25 40.12 18.67
CBDM (Qin et al., 2023) 24.12 35.02 48.07 20.90
OCLT (Zhang et al., 2024) 20.64 30.62 39.67 17.72
CCUA (ours) 21.32 27.98 30.39 15.24
Places-LT DDPM (Ho et al., 2020) 19.26 20.31 23.74 13.89
CBDM (Qin et al., 2023) 21.26 21.68 26.07 15.15
OCLT (Zhang et al., 2024) 19.16 19.43 22.84 13.04
CCUA (ours) 18.15 19.27 20.84 11.99
Results Across Various Category Intervals To see the effectiveness of our method on tail classes, we divide the Places-LT and TinyImageNet-LT datasets into three super categories: ‘Head’, ‘Body’, and ‘Tail’, with classes sorted by the number of training images. For each dataset, the top 33% classes were allocated to the ‘Head’ category, the next 34% classes to the ‘Body’ category, and the last 33% classes to the ‘Tail’ category. The percentage $P_{category}$ of training images belonging to each category is shown in the first row of Tab. 3. Our method achieves the best FID score across all three categories on Places-LT and TinyImageNet-LT datasets. For the ‘Tail’ classes, which are only $\sim 3\%$, our method improves their FID from 23.74 to 20.84 on Places-LT and 40.12 to 30.39 on TinyImageNet-LT, respectively. Additionally, both ‘Head’ and ‘Body’ classes show improved FID scores with our method, as demonstrated in Tab. 3.
4.3 Qualitative Results and Ablation Study
Visualization of Synthetic Images Fig. 4 shows synthetic images of SiT and with CCUA for ImageNet-LT tail classes ‘espresso’ and ‘window shade’. Compared to SiT, CCUA achieves visually higher fidelity and diversity. More qualitative results on ImageNet-LT, TinyImageNet-LT, and CIFAR100-LT are shown in appendix A.3.1 and A.3.3.
(a) Synthetic images from SiT (Ma et al., 2024).
(b) Synthetic images from CCUA (ours).
Figure 4: Synthetic images of SiT and it with CCUA for ImageNet-LT tail classes ‘espresso’ and ‘window shade’. All methods start the denoising process from the same Gaussian noise at corresponding grid cells. CCUA shows consistently higher diversity and fidelity compared to SiT.
Table 4: Loss Function Analysis.
[…]
$\mathcal{L}_{ddpm}$ 18.67
$\mathcal{L}_{ddpm}$ w/ $\mathcal{L}_{ucl}$ (Eq. 1) 18.05
$\mathcal{L}_{ddpm}$ w/ $\mathcal{L}_{al}$ (Eq. 6) 17.84
$\mathcal{L}_{ddpm}$ w/ $\mathcal{L}_{ccua}$ (Eq. 7) 17.16
Ablation Study of Loss Function Table 4 shows an ablation study of the proposed losses including $\mathcal{L}_{ucl}$ and $\mathcal{L}_{al}$, which is conducted on TinyImageNet-LT. For the loss function ablation study, we disable batch resample strategy which is separately ablated. Our method achieves lower FID when applying only the $\mathcal{L}_{ucl}$ loss (Eq. 1) or only $\mathcal{L}_{al}$ loss (Eq. 6) compared to the standard $\mathcal{L}_{ddpm}$ loss only. The best result is achieved with both losses, i.e., $\mathcal{L}_{ccua}$ (Eq. 7).
Ablation Study of Batch Resample StrategyAs shown in Table 5, we measure the performance of SiT and CCUA, with or without applying batch resampling strategy. Compared to SiT, the proposed CCUA loss achieves better IS, FID, sFID and Recall with only 0.09 Precision’s decline, illustrates our method on improving diversity without trading off fidelity. With applying batch resample strategy, the CCUA further improves all metrics. Notes that the CCUA increases Recall from 20.66 to 23.05, while SiT suffers a decline in Recall from 19.17 to 17.26 when applying batch resampling. As we discuss in Section. 3.1, the $\mathcal{L}_{ucl}$ could benefit from batch resample strategy due to the increasing instances of tail classes, leading to more diverse distribution in the latent space.
Table 5: Batch Resample Strategy Analysis on ImageNet-LT. All models are trained from scratch to 40 epochs. Blue ‘()’ highlights the improvement of each method compared to the SiT baseline, while Red ‘()’ highlights the decline compared to SiT baseline.
Method IS $\uparrow$ FID $\downarrow$ sFID $\downarrow$ Prec. (%) $\uparrow$ Recall (%) $\uparrow$
SiT (baseline) 53.93 33.87 22.66 54.56 19.17
  +Batch Resample 71.39 (+17.46) 28.05 (-5.82) 21.57 (-1.09) 57.61 (+3.05) 17.26 (-1.91)
$\mathcal{L}_{ccua}$ (Eq. 7) 57.79 (+3.86) 32.25 (-1.62) 20.71 (-1.95) 54.47 (-0.09) 20.66 (+1.49)
  +Batch Resample 73.60 (+19.67) 25.89 (-7.98) 16.54 (-6.12) 58.62 (+4.06) 23.05 (+3.88)
5 Conclusion
Real-world data for training image generation models often exhibit long-tailed distributions. Similar to class-imbalanced GANs (Khorram et al., 2024), class-imbalanced diffusion models generates inferior tail class images due to data scarcity. We propose a framework with two losses. Firstly, we employ an unsupervised contrastive loss with negative samples only to contrast the latents of different synthetic images at every denoising step, promoting intra-class diversity in particular for tail classes. Our second loss aligns class-conditional generation with unconditional generation for large timesteps. This encourages the initial denoising steps to be class-agnostic, thereby enriching tail classes through knowledge sharing from head classes–a principle demonstrated to enhance long-tailed GAN performance (Shahbazi et al., 2022; Khorram et al., 2024), which we successfully adapt to diffusion models. With the two losses, our framework of contrastive conditional-unconditional alignment boosts the performane of DDPM (Ho et al., 2020) and SiT (Ma et al., 2024) for long-tailed image generation and outperforms alternative methods including CBDM (Qin et al., 2023), OCLT (Zhang et al., 2024), and Dispersive Loss (Wang & He, 2025). Extensive experiments on multiple datasets in particular ImageNet-LT with 256x256 resolution demonstrated the effectiveness of our method on both UNet-based architecture and Diffusion Transformer.
Appendix A Technical Appendices and Supplementary Material
(a) w/ UNet-based Model
(b) w/ Diffusion Transformer
Figure 5: Model details of CCUA framework with UNet-based Model and Diffusion Transformer. As shown in Fig. 2, the noise estimation network is divided into two parts, a latent encoded network e()e(*) and a decoded network d()d(*). (a) For UNet-based architecture, e()e(*) is defined as the UNet encoder, while d()d(*) is defined as the UNet decoder. (b) For Diffusion Transformer, e()e(*) is defined as all the SiT/DiT blocks, while d()d(*) is defined as the final linear and adaLN projection layer.
Table 6: Our method outperforms other baselines on all datasets. We also provide the results of DDPM trained on balanced datasets, which show the upper bound of performance. All models are measured with DDIM (Song et al., 2020) 100 steps for conditional generation with CFG. Blue ‘()’ shows improvement of our method over DDPM baseline (Ho et al., 2020). Green ‘()’ shows improvement of DDPM trained on balanced version over trained on long-tailed version.